The length of a triangle is 3cm longer than its width. If the perimeter of the rectangle is 30cm, find its area.
Answer by JulietG(1812)

The perimeter is a fence around. P = 2L + 2W
You're asking first about a triangle and then about a rectangle. Let's solve for a rectangle, and you can repost if you need a different angle.
L = W + 3
Substitute the value of L into the first equation
P = 2L + 2W
30 = 2(W+3) + 2W
Distribute
30 = 2W+6 + 2W
30 = 4W + 6
Subtract 6 from each side
24 = 4W
Divide each side by 4
6 = W
If the width is 6, then the length is 9
Perimeter = 2(9) + 2(6)
30 = 18 + 12
30=30
Success!